# Decoherence and Probability

### v2

@@ 1. Introduction
These remarks prove enlightening when considered in the context of discussions of probability and emergence in the Many Worlds or Everett approach to the interpretation of quantum theory. In particular, consider the non-probabilistic emergentist account of Wallace (2012), Saunders (2021b) and Franklin (2023) in which features of the evolution equations of quantum theory are claimed to be sufficient, in some contexts to some degree, to justify a link between the Born weightings and physical salience, without these weightings being understood probabilistically. Decoherence, on this view, is a dynamical process under which probabilities emerge without the need for any prior probabilistic assumptions.111It should be noted that Franklin (2023) only explicitly claims to have established the non-probabilistic emergence of ‘classical structures’ rather than probability.
The non-probabilistic emergentist account of probability via decoherence is, however, unconvincing. One cannot justifiably presuppose the physical salience of structures derived via decoherence theory based upon an entirely uninterpreted use of the quantum formalism. Justifications of the physical salience of structures derived via decoherence theory purely based upon similarity to interpreted classical physics can never be satisfactory as an approach to the interpretation of physical theory. Some prior generalised concept of quantum measure or quantum probability as a weighting of possibilities must be assumed in the application of decoherence theory. Non-probabilistic emergentism about probability based on quantum decoherence fails. Nothing comes from nothing.
In this vein, it is worth noting that the probabilities which come into quantum theory via the experimentally confirmed expectation values cannot be represented as measures induced by the integration of a genuine probability density function over phase space. Rather, they can, at best, be represented in terms of the marginal probability distributions for position and momentum considered separately, with the density function taking the form of Wigner function, which is formally a quasi-probability density function in a quantum phase space representation (Wigner 1932, 1971).
@@
In what follows we will construct a novel account of the emergence of probability within the quantum phase formalism. Our approach takes as its starting point a partially interpreted quantum quasi-probability structure. We draw crucially upon the framing of emergence due to Butterfield (2011) and Palacios (2022) and of the semi-classical limit due to Feintzeig (2020). We show that the combined application of decoherence and semi-classical averaging, which neglects terms , leads to a classical probability model as an emergent coarse-grained description. This is not, of course, to offer a solution to the measurement problem in terms of a full interpretation of the relevant possibility spaces. Rather, what we offer is a conceptual framework for the analysis of classical and quantum probability within which any coherent interpretation must be expected to operate. Our aim is to clarify in which sense the principles of quantum mechanics require probabilistic concepts that could be connected to probabilistic measurement outcomes in a cogent way once a consistent interpretation of quantum mechanics has been established.
@@
The structure of the paper is as follows. Section 2 provides detailed exegesis and rebuttal of the non-probabilistic emergentist account of probability via decoherence focusing on the role of similarity arguments. Section 3 then provides a formal reconstruction of classical stochastic and quantum phase space mechanics to allow direct comparison as classical and quantum possibility space models. Section 4 demonstrates the sense in which classical possibility space models can be understood to emerge from quantum possibility space models. Decoherence alone is not sufficient to account for the emergence of classical probability from the underlying quantum dynamics. Rather, we require the combination of quasi-probabilistic emergence based on decoherence leading to a positive-definite quasi-probability function and coarse-grained emergence based upon a semi-classical averaging of observables which neglect terms . Section 5 recapitulates our results in the context of the motivating dialectic and offers some thoughts regarding outstanding issues of interest.
@@ 2.1. Decoherence and Emergence
The role of probability in the interpretation of quantum mechanics takes centre stage in the context of the relationship between the Everett interpretation and decoherence. In particular, according to what might be called the Oxford approach the Born rule can be extracted from the Many Worlds branching structure based on principles of reasoning that leave the application of the Born rule as the only rational way of betting on quantum outcomes open to an agent on an Everettian branch of the wave function who endorses the Everett interpretation.222See Deutsch (1999) and Wallace (2002) for the original proofs and Wallace (2012) for the systematic treatment in the context of the emergentist view. See also Saunders (2004, 2005); Wallace (2009); Greaves and Myrvold (2010). Various critical responses (and counter responses) are Price (2010); Rae (2009); Dizadji-Bahmani (2013); Adlam (2014); Dawid and Thébault (2014); Read (2018); Brown and Porath (2020); Steeger (2022); March (2023). For other approaches to probability in Many Worlds Theory see, for example, Saunders (2021a); Short (2023); Saunders (2024).
Zurek (2005) and Baker (2007) have criticized the Oxford approach by pointing at a circularity in its line of reasoning: decoherence must already be assumed to establish the branching of the wave function that provides the basis for identifying an agent who can consider betting along the lines of the decision theoretic argument. But decoherence already relies on a probabilistic interpretation of the process. In this context, Dawid and Thébault (2015) have argued that the the situation is even worse: the notion of probability required for understanding decoherence in the sense of a probabilistic suppression of off-diagonal elements of the density matrix is stronger than the decision theoretic notion of probability offered. This approach to extracting the Born rule therefore is not just circular but incoherent.
Against these criticisms, Saunders (2021b) and Franklin (2023) have sought to buttress the emergentist approach building on earlier discussions by Wallace (2010, 2012) and Saunders (2005). Franklin argues that “the neglect of terms with relatively small amplitudes can be justified non-probabilistically […] in contexts where interference is rife, the probabilistic interpretation of the (mod-squared) amplitudes is ruled out […] the Born rule, in such contexts, takes the form of an averaging measure rather than a probability measure. […] we should think of the relation between small amplitudes and irrelevance as a dynamical phenomenon. The relative magnitude of the amplitudes encodes the dynamical contribution of each term.”
@@
The original presentation of this form of similarity via dynamical irrelevance argument can be found within the highly influential emergentist defence and development of the Many Worlds interpretation due to Wallace (2012), who suggests that “[w]e can think of the significance of the Hilbert space metric as telling us when some emergent structure really is robustly present, and when it’s just a ‘trick of the light’ that goes away when we slightly perturb the microphysics…What makes perturbations that are small in Hilbert-space norm ‘slight’, [is] not the probability interpretation of them. Ultimately, the Hilbert- space norm is just a natural measure of state perturbations in Hilbert space, and that naturalness follows from considerations of the microphysical dynamics, independent of higher-level issues of probability” (pp. 253–254).
What Wallace, Franklin and Saunders all seem to have in mind is that features of the evolution equations of quantum theory are sufficient, in some contexts to some degree, to justify a link between the Born weightings and physical salience, without these weightings being understood probabilistically in any sense. The problems with such a strong emergentist view will be considered in the following section in the context of the reliance on uninterpreted similarity arguments.
@@ 2.2. Similarity and Interpretation
The failure of non-probabilistic emergence based on uninterpreted similarity arguments can be understood to arise from a basic conflict with the principle that a scientific theory should allow for empirical testing on its own terms. The key problem is the assumption that the set of rules that specify an important part of the theory’s empirical import, namely the decoherence of branches of the wave function, can be extracted from observing structural similarities to a theory that serves as a limiting case of that theory – the model where coherence terms are set to zero. In other words, a limiting theory serves as the basis for extracting empirical implications of the fundamental theory.
@@
The problem with this line of reasoning is that it does not explain what measuring a certain value of an observable implies at the level of the full theory. As long as no such understanding is forthcoming at the level of the full theory, however, we have no basis to decide whether or not we are justified to call any other theory a limiting theory of our full theory. Mere similarity arguments are insufficient for making that decision for one principal reason: as long as the implications of measurements cannot be spelled out at the level of the full theory, we remain insensitive to the distinction between empirically relevant stable dynamics on the one hand and spurious dynamics of parameterization prescriptions on the other. In the limiting theory that sets coherence effects to zero, the set of allowed states are confined to states that show no coherence effects. Any discovery of coherence effects would therefore contradict the limiting theory. The question as to whether coherence should be considered probable or improbable thus does not arise. Coherence is ruled out. In the full theory, coherence is consistent with the theory. Coherence effects are represented in the theory’s set of allowed states. To understand whether they are suppressed or not, it is not sufficient to point at a small dimensionless number that characterizes cohered states because small dimensionless numbers might, in principle, also be extracted from specific parameterizations of the theory that bear no physical significance. To rule out this possibility, one needs to find the basis for a probabilistic analysis of those states at the level of the full theory.
@@
The similarity approach has a second, related problem. While a limiting theory can be deduced from a fundamental theory, the opposite is not true. A probabilistic interpretation of the non-cohered limiting theory (to the extent it can be given) does not formally imply the probabilistic characteristics of the fully quantum regime in terms of the full Born rule. In other words, we end up deploying two entirely different lines of reasoning to establish what formally looks like one coherent concept of quantum probability. All this is a far cry from the initial claim that Many Worlds quantum mechanics has the attractive feature to require no posits beyond the wave function equations. Indeed, an appeal to decoherence as a precondition of interpretational content would render the Many Worlds approach of a piece with precisely the pragmatic, neo-Bohrian outlook that the Many Worlds view motivated by rejecting. For example, such an approach would involve implementing the proscription on the use of the Born rule as a probabilistic rule due to Healey (2017). Pragmatic approaches to quantum theory are without doubt interesting and valuable in their own rights. However, we do not take a marriage with the Many Worlds view of quantum mechanics to be a union that would be to the profit of either party.
@@
In summary, returning to our principal argument: As long as no probabilistic interpretation of the wave function is provided at the level of quantum mechanics, it is not clear whether Born weights are a characteristic of physically relevant dynamics or of mere parameterization. Therefore, it is not justified to infer the empirical import of Many Worlds quantum mechanics from the fact that the resulting wave function in a given limit looks strikingly similar to the empirical results of a non-cohered theory. One might assert by fiat that the import of Many Worlds quantum mechanics matches the import of the corresponding non-cohered theory in a given limit. If one goes down that road, however, the non-cohered theory turns from a limiting theory of the full quantum theory into an essential element of quantum theory that is needed for providing the link between the theory’s formal structure and its empirical import. The result is a confusing compound of mutually dependent theoretical posits. We cannot make valid inferences about the world based upon uninterpreted similarity arguments combined with the formal structure of a decoherence model.
@@ 3. Probability and Possibility
Franklin thus argues that it is misguided to even look for a probabilistic interpretation of the dynamics at the quantum level because quantum theory provides no basis for a quantum probability measure that satisfies the Kolmogorov axioms.333There are well-known problems with constructing probability measures over the full set of states of a quantum theory. In particular, not only are there general grounds for thinking quantum interference is in tension with any probabilistic interpretation of the quantum amplitudes (Wallace 2014), but one can in fact show that there is a close formal relation between the absence of a well-defined joint probability distribution for non-commuting observables and violation of the Bell-CHSH inequalities (Fine 1982a, b; Pitowsky 1989; Suppes and Zanotti 1993; Hartmann 2015). On his reasoning, Zurek (2005), Baker (2007) and Dawid and Thébault (2015) are not just wrong in claiming that decoherence needs to be based on a probabilistic interpretation of quantum processes. They are already wrong in assuming that a probabilistic interpretation of the quantum regime is a meaningful goal. Franklin asserts that establishing a probabilistic account at the level of the limiting theory is the only way to get from quantum mechanics to empirical predictions. In this section and the next we will carry out a detailed analysis of this issue and present a proposal for the precise sense in which probabilistic concepts can be identified at the level of the full theory.
@@ 3.1.1. Quasi-Probability Structures
A quasi-probability structure is a triple where the three elements are defined as follows:444Here we are using a slight generalisation of the framework set out in Dowker and Wilkes (2022).
@@
Event Algebra: is a non-empty collection of sub-sets of such that:
@@
By definition we have that , , is closed under-finite intersection, and
Two important features that a quasi-probability model need not have are -additivity and positivity. The first is since we have not insisted that the event algebra is a -algebra; it need not be closed under countable unions.555For a detailed discussion of relationship between forms of additivity and classical and quantum probabilities see Arageorgis et al. (2017). The second is since we have not insisted that the quasi-measure is a measure; it need not be positive (nor indeed -additive). Strengthening the model to include both these features results in the familiar formal structure of a classical probability model.
@@ 3.1.2. Classical Probability Structures
-Algebra: is a non-empty collection of sub-sets of such that:
@@
for all (closed under comeplementation);
@@ 3.2. Partial Interpretation via Possibility Weightings
The concepts ‘possible’, ‘impossible’, and ‘certain’ can be given their natural linguistic meaning. The crucial concept of ‘higher possibility weight’, by contrast, requires some further articulation to be fully understood. We do not intend to provide this here, but rather articulate the framework in which any such articulation must work. Clearly, an event being higher weighted than another event indicates that it is quantitatively closer to certain events and further from impossible events, with the closeness given by the quasi-measure. However, the concept of higher possibility weighting is a only a partial interpretation of the quasi-measure structure in that it is an assignment of meaning to a structure which allows for a range of further admissible interpretation. In the case of classical probability and measures, at least, such further admissible interpretations can be directly provided in terms of standard notions of credence, frequency, or physical chance. In the case of a quasi-probability the structure of admissible interpretations are required to be more exotic or include further ingredients. We will return to this issue in our closing remarks.

### v3

@@ 1. Introduction
These remarks prove enlightening when considered in the context of discussions of probability and emergence in the Many Worlds or Everett approach to the interpretation of quantum theory. In particular, consider the non-probabilistic emergentist account of Wallace (2012), Saunders (2021b) and Franklin (2023) in which features of the uninterpreted formalism of quantum theory are claimed to be sufficient, in some contexts to some degree, to justify a link between the Born weightings and physical salience. Decoherence, on this view, is designated a dynamical process under which stable, quasi-classical, ‘branching structure’ emerges. This emergence is intended to be entirely independent of probabilistic interpretive assumptions. The link between the Born weightings and physical salience is established either directly within the quantum formalism via ‘structural stability’ arguments or indirectly based upon similarity between the quantum formalism and interpreted classical physics.
In Section 2 we argued towards the failure of this approach. First, one cannot justifiably presuppose the physical salience of structures derived via decoherence theory based upon an entirely uninterpreted use of the quantum formalism. Second, one cannot coherently utilise appeal to the limiting classical theory as a requirement for interpretational content, and assert that one is presenting an interpretation of the theory at a quantum level, as opposed to a pragmatic prescription for its application in decohered contexts. Rather, in the context of any putative interpretation of quantum theory, a generalised concept of quantum measure or quantum probability as a weighting of possibilities must be assumed in the application of decoherence theory. The non-probabilistic emergentist approach to quantum decoherence fails since it denies itself precisely the conceptual resources needed to link Born weightings with physical salience. Nothing comes from nothing.
@@
In what follows we will construct a novel account of the emergence of probability within the quantum phase space formalism. Our approach takes as its starting point a partially interpreted quantum quasi-probability structure within the quantum phase space approach (Wigner 1932, 1971). We draw crucially upon the framing of emergence due to Butterfield (2011) and Palacios (2022) and of the semi-classical limit due to Feintzeig (2020). We show that the combined application of decoherence and semi-classical averaging, which neglects terms , leads to a classical probability model as an emergent coarse-grained description. This is not, of course, to offer a solution to the measurement problem in terms of a full interpretation of the relevant possibility spaces. Rather, what we offer is a conceptual framework for the analysis of classical and quantum probability within which any coherent interpretation of quantum mechanics must be expected to operate. Our aim is to clarify in which sense the principles of quantum mechanics require probabilistic concepts that could be connected to probabilistic measurement outcomes in a cogent way once a consistent interpretation of quantum mechanics has been established.
@@
The structure of the paper is as follows. Section 2 provides detailed exegesis and rebuttal of the non-probabilistic emergentist account of probability via decoherence focusing on the role of similarity arguments. Section 3 then provides a formal reconstruction of classical stochastic and quantum phase space mechanics to allow direct comparison as classical and quantum possibility space models. Section 4 demonstrates the sense in which classical possibility space models can be understood to emerge from quantum possibility space models based upon the combination of quasi-probabilistic emergence via decoherence leading to a positive-definite quasi-probability function and coarse-grained emergence via semi-classical averaging of observables which neglect terms . Section 5 recapitulates our results in the context of the motivating dialectic and offers some thoughts regarding outstanding issues of interest.
@@ 2.1. Decoherence and Emergence
The role of probability in the interpretation of quantum mechanics takes centre stage in the context of the relationship between the Everett interpretation and decoherence. In particular, according to what might be called the Oxford approach the Born rule can be extracted from the Many Worlds branching structure based on principles of reasoning that leave the application of the Born rule as the only rational way of betting on quantum outcomes open to an agent on an Everettian branch of the wave function who endorses the Everett interpretation.111See Deutsch (1999); Wallace (2002, 2012); Saunders (2004, 2005); Wallace (2009); Greaves and Myrvold (2010); Price (2010); Rae (2009); Dizadji-Bahmani (2013); Adlam (2014); Dawid and Thébault (2014); Read (2018); Brown and Porath (2020); Steeger (2022); March (2023); Saunders (2021a); Short (2023); Saunders (2024).
Zurek (2005) and Baker (2007) point to a circularity in Oxford line of reasoning: decoherence must already be assumed to establish the branching of the wave function that provides the basis for identifying an agent who can consider betting along the lines of the decision theoretic argument. But decoherence already relies on a probabilistic interpretation of the process. In this context, Dawid and Thébault (2015) diagnose a deeper problem: the notion of probability argued to be required for understanding decoherence in the sense of a probabilistic suppression of off-diagonal elements of the density matrix is taken to be stronger than the decision theoretic notion of probability offered. This approach to extracting the Born rule therefore is not just circular but incoherent.
The original non-probability presentation of decoherence in the Oxford view is due to Wallace (2010, 2012), who suggests that “[w]e can think of the significance of the Hilbert space metric as telling us when some emergent structure really is robustly present, and when it’s just a ‘trick of the light’ that goes away when we slightly perturb the microphysics…What makes perturbations that are small in Hilbert-space norm ‘slight’, [is] not the probability interpretation of them. Ultimately, the Hilbert- space norm is just a natural measure of state perturbations in Hilbert space, and that naturalness follows from considerations of the microphysical dynamics, independent of higher-level issues of probability” (pp. 253–254). However, as noted by Dawid and Thébault (2015), this purely structural emergentist account provides no justification for the connection between Born weightings and physical salience.
In this context, Saunders (2021b) and Franklin (2023) have sought to further develop Wallace’s approach. Franklin argues that “the neglect of terms with relatively small amplitudes can be justified non-probabilistically […] in contexts where interference is rife, the probabilistic interpretation of the (mod-squared) amplitudes is ruled out […] the Born rule, in such contexts, takes the form of an averaging measure rather than a probability measure. […] we should think of the relation between small amplitudes and irrelevance as a dynamical phenomenon. The relative magnitude of the amplitudes encodes the dynamical contribution of each term.”
@@
What Wallace, Franklin and Saunders all seem to have in mind is that features of the evolution equations of quantum theory are sufficient, in some contexts to some degree, to justify a link between the Born weightings and physical salience, without these weightings being understood probabilistically in any sense. The justification for this connection is then provided in terms of structural similarity to the classical theory. The problems with such a emergentist view will be considered in the following section in the context of the reliance on uninterpreted similarity arguments.
@@ 2.2. Similarity and Interpretation
The failure of non-probabilistic emergence based on uninterpreted similarity arguments can be understood to arise from a basic conflict with the principle that a scientific theory should allow for empirical testing on its own terms. The key problem is the assumption that the set of rules that specify an important part of the theory’s empirical import, namely the decoherence of branches of the wave function, can be extracted from observing structural similarities to a theory that serves as a limiting case of that theory – the model where coherence terms are set to zero. In other words, a ‘limiting theory’ serves as the basis for extracting empirical implications of the fundamental theory.222Here and below we assume the ‘factual’ understanding limits as per Section §4.3.
@@
The problem with this line of reasoning is that it does not explain what measuring a certain value of an observable means at the level of the full theory. As long as no such understanding is forthcoming at the level of the full theory, however, we have no basis to decide whether or not we are justified to call any other theory a limiting theory of our full theory. Mere similarity arguments are insufficient for making that decision for one principal reason: as long as the implications of measurements cannot be spelled out at the level of the full theory, we remain insensitive to the distinction between empirically relevant stable dynamics on the one hand and spurious dynamics of parameterization prescriptions on the other. In the limiting theory that sets coherence effects to zero, the set of allowed states are confined to states that show no coherence effects. Any discovery of coherence effects would therefore contradict the limiting theory. The question as to whether coherence should be considered probable or improbable thus does not arise. Coherence is ruled out. In the full theory, coherence is consistent with the theory. Coherence effects are represented in the theory’s set of allowed states. To understand whether they are suppressed or not, it is not sufficient to point at a small dimensionless number that characterizes cohered states because small dimensionless numbers might, in principle, also be extracted from specific parameterizations of the theory that bear no physical significance. To rule out this possibility, one needs to find the basis for those states at the level of the full theory. Such an analysis requires appeal to a generalised concept of quantum measure or quantum probability as a weighting of possibilities.
@@
The similarity approach has a second, related problem. While a limiting theory can be deduced from a fundamental theory, the opposite is not true. A probabilistic interpretation of the limiting theory with zero coherence (to the extent it can be given) does not formally imply the probabilistic characteristics of the fully quantum regime in terms of the full Born rule. In other words, we end up deploying two entirely different lines of reasoning to establish what formally looks like one coherent concept of quantum probability. All this is a far cry from the initial claim that Many Worlds quantum mechanics has the attractive feature to require no posits beyond the wave function equations. Indeed, an appeal to decoherence as a precondition of any interpretational content would render the Many Worlds approach of a piece with precisely the pragmatic, neo-Bohrian outlook that is inconsistent for any claim of realism about the quantum state, whether presented in functionalist/structuralist terms or otherwise.333See (Wallace 2012, §1.8). What we say here is perfectly compatible with the (very reasonable) view that the ‘worlds’ in Many Worlds require decoherence to be attributed interpretational content. For example, such an approach would involve implementing the precisely the prescription on the use of the Born rule as a probabilistic rule advocated by Healey (2017). Pragmatic approaches to quantum theory are without doubt interesting and valuable in their own rights. However, we do not take a marriage with the Many Worlds view of quantum mechanics to be a union that would be to the profit of either party.
@@
In summary, returning to our principal argument: As long as no probabilistic interpretation of the wave function is provided at the level of quantum mechanics, it is not clear whether Born weights are a characteristic of physically relevant dynamics or of a spurious parameterization. Therefore, it is not justified to infer the empirical import of Many Worlds quantum mechanics from the fact that the resulting wave function in a given limit looks strikingly similar to the empirical results of a zero coherence theory. One might assert by fiat that the import of Many Worlds quantum mechanics matches the import of the corresponding zero coherence theory in a given limit. If one goes down that road, however, the zero coherence theory turns from a limiting theory of the full quantum theory into an essential element of quantum theory that is needed for providing the link between the theory’s formal structure and its empirical import. The result is a confusing compound of mutually dependent theoretical posits. We cannot make valid inferences about the world based upon uninterpreted similarity arguments combined with the formal structure of a decoherence model.
@@ 3. Probability and Possibility
Franklin thus argues that it is misguided to even look for a probabilistic interpretation of the dynamics at the quantum level because quantum theory provides no basis for a quantum probability measure that satisfies the Kolmogorov axioms.444See (Fine 1982a, b; Pitowsky 1989; Suppes and Zanotti 1993; Hartmann 2015; Wallace 2014). On his reasoning, Zurek (2005), Baker (2007) and Dawid and Thébault (2015) are not just wrong in claiming that decoherence needs to be based on a probabilistic interpretation of quantum processes. They are already wrong in assuming that a probabilistic interpretation of the quantum regime is a meaningful goal. Franklin asserts that establishing a probabilistic account at the level of the limiting theory is the only way to get from quantum mechanics to empirical predictions. In this section and the next we will carry out a detailed analysis of this issue and present a proposal for the precise sense in which probabilistic concepts can be identified at the level of the full theory.
@@ 3.1.1. Quasi-Probability Structures
A quasi-probability structure is a triple where the three elements are defined as follows:
@@
Event Algebra: is a non-empty collection of subsets of such that:
@@
By definition we have that , , is closed under finite intersection, and
Two important features that a quasi-probability model need not have are -additivity and positivity. The first is since we have not insisted that the event algebra is a -algebra; it need not be closed under countable unions, cf. (Arageorgis et al. 2017). The second is since we have not insisted that the quasi-measure is a measure; it need not be positive (nor indeed -additive). Strengthening the model to include both these features results in the familiar formal structure of a classical probability model.
@@ 3.1.2. Classical Probability Structures
-Algebra: is a non-empty collection of subsets of such that:
@@
for all (closed under complementation);
@@ 3.2. Partial Interpretation via Possibility Weightings
The concepts ‘possible’, ‘impossible’, and ‘certain’ can be given their natural linguistic meaning. The crucial concept of ‘higher possibility weight’, by contrast, requires some further interpretation to be fully understood. We do not intend to provide this here, but rather set-out the framework in which any such interpretation must work. An event being higher weighted than another event indicates that it is quantitatively closer to certain events and further from impossible events, with the closeness given by the quasi-measure. However, the concept of higher possibility weighting is a only a partial interpretation of the quasi-measure structure in that it is an assignment of meaning to a structure which allows for a range of further admissible interpretation. In the case of classical probability and measures, such further admissible interpretations can be directly provided in terms of standard notions of credence, frequency, or physical chance. In the case of a quasi-probability the structure of admissible interpretations are required to be more exotic or include further ingredients. We will return to this issue in our closing remarks.